A councilwoman regularly polls her constituency to gauge her level of support among voters. Her support...

  1. A councilwoman regularly polls her constituency to gauge her level of support among voters. Her support is usually at 50% among his constituents. If a sample of 100 people is selected answer the following questions:
    1. What is the mean of sampling distribution for sample proportions?
    2. What is the standard error for sample for sample proportions?
    3. What is the probability that the councilwoman support falls below 45%?
Answer #1

Solution

Given that,

The mean of the sampling distribution of proportion is ,

= p = 0.50

The standard  standard error for sample for sample proportions is,

=  [p( 1 - p ) / n = [(0.50 * 0.50) / 100 ] = 0.05

(c)

P( < 0.45) =

= P[( - ) / < (0.45 - 0.50) / 0..05]

= P(z < -1)

= 0.8413